Currently, two galaxies similar to the Milky Way and Andromeda (M31) galaxies are separated by about 2.70 million light-years and moving toward each other at about 115.00 km/s. Estimate how long it may take for the two to collide.

____ years

Compared to the speed of light, rank the following astronomical phenomena from slowest to fastest speed.

  • Jet from an AGN

  • X-ray light

  • Fastest cosmic ray

  • Solar wind
Answer #1

A)

given
v = 115 km/s = 1.15*10^5 m/s
d = 2.7 millian light year

= 2.7*10^6*3*10^8*365*24*60*60 m

= 2.5544*10^22 m

time taken for them to meet,

t = distance between galaxies/relative speed of galaaxies

= d/(2*v)

= 2.5544*10^22/(2*1.15*10^5)

= 1.11*10^17 s

= 1.11*10^17/(365*24*60*60) year

= 3.52*10^9 years <<<<<<<<<<------------Answer

B)
X-ray light > Fastest cosmic ray > Solar wind > Jet from an AGN